OpenStudy (anonymous):

i know the answer is not A but i think it is D The sum of two numbers is 110. Their difference is 28. What is the smallest of the two numbers? A. 56 B. 41 C. 69 D. 55

OpenStudy (mani_jha):

Its B. 41. The two nos are 41 and 69

OpenStudy (anonymous):

(110-28)/2 = ?

OpenStudy (anonymous):

i think it one is a but not sure For which system of equations would elimination be the simplest technique to choose? A. y = 2x and x + 3y = 5 B. 4x + 3y = 6 and 3y = 5 C. x + y = 7 and x - y = 12 D. 3x - 2y = 8 and 4x - 5y = 20

OpenStudy (anonymous):

i think it is a